dinitrogen difluoride gas is collected at 5.o c in an evacuated flask with a mesuared volume of 50.0 L. When all gas has been collected, the pressure in the flask is measured to be 0.090 atm
calculate the mass and number of moles of dinitrogen difluoride gas that were collected. round to 2 significant digts
1
Expert's answer
2019-07-25T05:30:20-0400

1)Use Ideal Gas Law to find moles of a gas:

PV =nRT

P = 0.090 atm

V = 50.0 L

T(K)= 5+273= 278 K

R = 0.082 L*mol/atm*K


0.090*50.0 = n*0.082*278

n = 0.20 mol


2) find mass os gas:

m = M*n

M(N2F2) = 0.2*66.01 = 13 g


Answer: 1) 0.20 mol, 2) 13 g